OpenStudy (anonymous):

if x+iy=(1- i√3)^100,find x,y.

OpenStudy (anonymous):

\[x+iy=(1- i√3)^{100}\]

OpenStudy (kc_kennylau):

Express \(1-i\sqrt3\) in the form of \(r(\cos\theta+i\sin\theta)\) first

OpenStudy (anonymous):

how?

OpenStudy (kc_kennylau):

Any complex number in the form of \(a+bi\) can be expressed in the form of \(r(\cos\theta+i\sin\theta)\), where \(r=\sqrt{a^2+b^2}\) and \(\theta=\tan^{-1}\dfrac ba\)

OpenStudy (kc_kennylau):

For example, \(6+2i=\sqrt{6^2+2^2}(\cos(\tan^{-1}\dfrac26)+i\sin(\tan^{-1}\dfrac26))\)

OpenStudy (anonymous):

I get : 2^99(1-√3i)^100

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

You forgot to change the things inside to cosine and sine?

OpenStudy (anonymous):

no no, after solving cos and sin i get that..

OpenStudy (kc_kennylau):

For example, \(1-i=\sqrt{1^2+1^2}(\cos(\tan^{-1}\frac{-1}1)+\sin(\tan^{-1}\frac{-1}1))=\sqrt2(\cos135^\circ+i\sin135^\circ)\)

OpenStudy (kc_kennylau):

Don't solve the cosine and sine

OpenStudy (anonymous):

@kc_kennylau : don't give examples please? Here, r=2, and theta= -pi/3

OpenStudy (anonymous):

there you get: 2(cos(-pi/3)+sin(-pi/3)) = 2(1/2-√3/2)

OpenStudy (anonymous):

sorry, = [2(1/2 - √3/2)] ^100

OpenStudy (anonymous):

= 2^99 (1-√3)^100 = 2^99(x+√3i)

OpenStudy (kc_kennylau):

So \((1-\sqrt3)^{100}=\{2[\cos(-\dfrac\pi3)+\sin(-\dfrac\pi3)]\}^{100}\)

OpenStudy (kc_kennylau):

Don't solve the cosine and sine

OpenStudy (kc_kennylau):

Or you'd have turned them into cosine and sine and then turned it back

OpenStudy (kc_kennylau):

Correction: \((1-\sqrt3)^{100}=\{2[\cos(-\dfrac\pi3)+i\sin(-\dfrac\pi3)]\}^{100}\)

OpenStudy (anonymous):

then what do i do?

OpenStudy (kc_kennylau):

\[=2^{100}[\cos(-\frac\pi3)+\sin(-\frac\pi3)]^{100}\]

OpenStudy (kc_kennylau):

Theorem: \(\displaystyle\Large(\cos\theta+i\sin\theta)^n=\cos(n\theta)+i\sin(n\theta)\)

OpenStudy (kc_kennylau):

Correction: \(\displaystyle=2^{100}[\cos(-\frac\pi3)+i\sin(-\frac\pi3)]^{100}\)

OpenStudy (kc_kennylau):

Apply this theorem

OpenStudy (kc_kennylau):

\[\begin{array}{rcl} (1-\sqrt3)^{100}&=&\{2[\cos(-\dfrac\pi3)+i\sin(-\dfrac\pi3)]\}^{100}\\ &=&2^{100}[\cos(-\frac\pi3)+i\sin(-\frac\pi3)]^{100}\\ &=&2^{100}[\cos(-\frac{100\pi}3)+i\sin(-\frac{100\pi}3)]\\ &=&2^{100}[\cos(-\frac{4\pi}3)+i\sin(-\frac{4\pi}3)]\\ &=&\mbox{now solve the cosine and sine} \end{array}\]

OpenStudy (anonymous):

AWESOME. I get : 2^99(-1 + √3i) == (which is the answer) :D

OpenStudy (kc_kennylau):

Glad you got your answer :)

OpenStudy (kc_kennylau):

Are you sure that you understood every step?

OpenStudy (anonymous):

Yep :)